what is the answeat to log8/logx=3
Answer by stanbon(48535)

what is the answer to log8/logx=3
-------
Cross-multiply to get:
3log(x) = log(8)
log(x^3) = log(8)
x^3 = 8
x = 2
===========
Cheers,
Stan H.